suppose that women's body temperatures are normally distributed with a mean of 98.4 F and a standard devidation of 0.70 F. what is the 95th percentile of womens body temperatures.
Answer by stanbon(75887)

suppose that women's body temperatures are normally distributed with a mean of 98.4 F and a standard devidation of 0.70 F. what is the 95th percentile of womens body temperatures
----------------
Find the z-value with a left-tail of 0.95:: invNorm(0.95) = 1.645
-----
Find the corresponding body temperature
-----
x = z*s + u
x = 1.645*0.7 + 98.4 = 99.55 degrees
----------------------------------
Cheers,
Stan H.
================